# A Free-fermion Formulation of Two-dimensional Ising Models

### v1

@@ III.3.2 Imaginary field
$Z=i^{N}\sum\limits_{\left\{{\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)}\right\}=\pm 1}{\omega\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)}$ (45)
@@
$\omega\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)=\sum\limits_{s_{5}=\pm 1}{(s_{1}s_{2}s_{5})e^{-\beta J{s_{5}}\left({s_{1}}+{s_{2}}+{s_{3}}+{s_{4}}\right)}\times}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}e^{-\beta J\left(s_{1}s_{2}+s_{3}s_{4}\right)}.$ (46)
@@
$\omega_{1}=e^{-6\beta J}-e^{2\beta J},$
@@
$\omega_{3}=\omega_{4}=\omega_{5}=\omega_{6}=\omega_{7}=\omega_{8}=0,$
$\omega_{9}=\omega_{11}=\omega_{13}=\omega_{15}=e^{2\beta J}-e^{-2\beta J},$
$\omega_{10}=\omega_{12}=\omega_{14}=\omega_{16}=-\omega_{9}.$ (47)
@@
$\tilde{\omega}_{9}=\frac{1}{2}e^{2\beta J}\left(3+e^{-4\beta J}\right)\left(1-e^{-4\beta J}\right),$
$\tilde{\omega}_{10}=-\frac{1}{2}e^{2\beta J}\left(1-e^{-4\beta J}\right)^{2},$
$\tilde{\omega}_{11}=\tilde{\omega}_{13}=\tilde{\omega}_{15}=\tilde{\omega}_{9},$
$\tilde{\omega}_{12}=\tilde{\omega}_{14}=\tilde{\omega}_{16}=\tilde{\omega}_{10}.$ (48)
It is trivial to see that this case is equivalent with an odd free-fermion model, the exact result is thereby given by the odd free-fermion solution in Eqs. (36) and (37)
@@
$=i\frac{\pi}{2}+\frac{1}{6}\ln\left[{\left({1+3{e^{4\beta J}}}\right){{\left|{1-{e^{-4\beta J}}}\right|}^{3}}}\right].$ (49)
The expression is surprisingly simple. To our knowlegde, this result has not been reported previously. There have been research works studying the Kagomé lattice Ising model in a magnetic field by transformation into the honeycomb lattice model [90, 91, 92]. One can examine that, the case in the imaginary field was not included in these works. Our analysis demonstrates that this case is also exactly solvable.
As is well-known, most of the exact solutions of two-dimensional Ising models are of the form of a double integral. Onsager’s solution of the square lattice model [5] is a very early example. Each model studied in this paper using the free-fermion approach involves a double integral in the solution, except for the Kagomé lattice model in the imaginary field. It is interesting to compare the result Eq. (49) with Eq. (1) of Ref. [93], which is the solution of the dimer model on the Kagomé lattice and can also be derived by mapping to the odd free-fermion model. Both formulas take a very simply logarithmic form. Ref. [93] stated that the novel and unique expression “points to the special role played by the Kagomé lattice”. Here we provide new evidence for this statement.
When $J>0$, the Kagomé lattice Ising model is geometrically frustrated. The residual entropy in the imaginary field is easily calculated from Eq. (49) by taking the zero temperature limit
@@ IV Summary
Each solution in this paper is achieved by mapping the Ising spin system into a free-fermion model. The mapping procedure is simple and easy to extend to other lattice systems. Applying the mapping method to the Ising model in a physical (real) and non-zero magnetic field would be interesting and likely very fruitful. We expect that our work provides new insights into the mathematical exact results of spin systems in a magnetic field. Furthermore, the free-fermion formulation may be suitable for other statistical lattice problems, such as the dimer covering problem. Study of the two-dimensional dimer models using the free-fermion approach is currently in progress.
@@ Acknowledgements.
This work was supported by Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2200001), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2021A1515010328) and National Natural Science Foundation of China (Grant No. 11874312).

### v2

@@ III.3.2 Imaginary-field case
$Z=i^{N}\sum\limits_{\left\{{\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)}\right\}=\pm 1}~{\prod\limits_{{\rm{all~vertices}}}{\omega\left(s_{1},s_{2},s_{3},s_{4}\right)}}$ (61)
@@
$\omega\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)=\sum\limits_{s_{5}=\pm 1}(s_{1}s_{2}s_{5})e^{-\beta{J_{3}}\left({{s_{1}}{s_{2}}+{s_{3}}{s_{4}}}\right)}\times$
$~~~~~~~~~~~~~~~~~~~~~~~~~~e^{-\beta\left[{{J_{1}}{s_{5}}\left({{s_{2}}+{s_{3}}}\right)+{J_{2}}{s_{5}}\left({{s_{1}}+{s_{4}}}\right)}\right]}.$ (62)
@@
${\omega_{1}}={e^{-2\beta{J_{3}}}}\left[e^{-2\beta\left(J_{1}+J_{2}\right)}-e^{2\beta\left(J_{1}+J_{2}\right)}\right],$
@@
${\omega_{3}}={e^{2\beta{J_{3}}}}\left[e^{-2\beta\left(J_{1}-J_{2}\right)}-e^{2\beta\left(J_{1}-J_{2}\right)}\right],$
$\omega_{4}=-\omega_{3},$
${\omega_{5}}={\omega_{6}}={\omega_{7}}={\omega_{8}}=0,$
$\omega_{9}=\omega_{11}=e^{2\beta J_{2}}-e^{-2\beta J_{2}},$
$\omega_{10}=\omega_{12}=-\omega_{9},$
$\omega_{13}=\omega_{15}=e^{2\beta J_{1}}-e^{-2\beta J_{1}},$
$\omega_{14}=\omega_{16}=-\omega_{13}.$ (63)
@@
$\tilde{\omega}_{9}={e^{-2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]-{e^{2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~+2\sinh(2\beta J_{2}),$
$\tilde{\omega}_{10}={e^{-2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]-{e^{2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~~-2\sinh(2\beta J_{2}),$
$\tilde{\omega}_{11}=\tilde{\omega}_{9},$
$\tilde{\omega}_{12}=\tilde{\omega}_{10},$
$\tilde{\omega}_{13}={e^{-2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]+{e^{2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~~+2\sinh(2\beta J_{1}),$
$\tilde{\omega}_{14}={e^{-2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]+{e^{2\beta{J_{3}}}}\sinh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~~-2\sinh(2\beta J_{1}),$
$\tilde{\omega}_{15}=\tilde{\omega}_{13},$
$\tilde{\omega}_{16}=\tilde{\omega}_{14}.$ (64)
It is straightforward to verify that ${\tilde{\omega}}_{9}{\tilde{\omega}}_{10}={\tilde{\omega}}_{13}{\tilde{\omega}}_{14}$ such that this case is equivalent with an odd free-fermion model, the exact result is thereby given by the odd free-fermion solution in Eqs. (52) and (53)
@@
[…]
$~~~~~~~~~~~~~~~~~~~~~~\left.\left.-2S_{1}S_{2}S_{3}C_{1}C_{2}C_{3}\right|\right]$ (65)
with $C_{i}=\cosh(2\beta J_{i})$ and $S_{i}=\sinh(2\beta J_{i})$ ($i=1,2,3$). The expression is surprisingly simple. To our knowlegde, this result has not been reported previously. There have been research works studying the Kagomé lattice Ising model in a magnetic field by transformation into the honeycomb lattice model Giacomini (1988a, b); Lu and Wu (2005). One can examine that, the case in the imaginary field was not included in these works. Our analysis demonstrates that this case is also exactly solvable.
We note that, the details of the mapping and calculation procedure are very important. In this case, the vertex weights are defined in Eq. (62), in which the factor $\prod\limits_{i=1}^{N}{{s_{i}}}$ in the partition function exhibits a subfactor ${s_{1}}{s_{2}}{s_{5}}$. One can see that, the effect of the field is partitioned as $e^{\beta{H_{{\rm{ex}}}}\left({{s_{1}}+{s_{2}}+{s_{5}}}\right)}$} for this vertex unit. Now we repartition this effect as $e^{\beta{H_{{\rm{ex}}}}\left[{\frac{1}{2}\left({{s_{1}}+{s_{2}}+{s_{3}}+{s_{4}}}\right)+{s_{5}}}\right]}$}, i.e., the vertex weights are reset as
$\omega\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)=\sum\limits_{{s_{5}}=\pm 1}\left(s_{1}^{1/2}s_{2}^{1/2}s_{3}^{1/2}s_{4}^{1/2}{s_{5}}\right)\times$
$~~~~~~~~~~~~~~~~{e^{-\beta{J_{3}}\left({{s_{1}}{s_{2}}+{s_{3}}{s_{4}}}\right)}}{e^{-\beta\left[{{J_{1}}{s_{5}}\left({{s_{2}}+{s_{3}}}\right)+{J_{2}}{s_{5}}\left({{s_{1}}+{s_{4}}}\right)}\right]}}$
with $1^{1/2}=1$ and ${(-1)}^{1/2}=i$. In this way, we will find that
${\omega_{1}}=-{\omega_{2}},{\rm{}}{\omega_{3}}=-{\omega_{4}},{\rm{}}{\omega_{5}}={\omega_{6}}={\omega_{7}}={\omega_{8}}=0,$
${\omega_{9}}={\omega_{10}},{\rm{}}{\omega_{11}}={\omega_{12}},{\rm{}}{\omega_{13}}={\omega_{14}},{\rm{}}{\omega_{15}}={\omega_{16}}.$
We obtain neither a symmetric nor an antisymmetric sixteen-vertex model. Hence, we fail to map the Ising model into an even or odd free-fermion model by weak-graph expansion. We will probably miss this new solvable case, or at least are not able to solve this case so conveniently by mapping into the odd free-fermion model as we have shown. This fact indicates that, the details of mapping and calculation may play the key role in leading to the new solution.
As is well-known, most of the exact solutions of two-dimensional Ising models are of the form of a double integral. Onsager’s solution of the square lattice model Onsager (1944) is a very early example. Each model studied in this paper using the free-fermion approach involves a double integral in the solution, except for the Kagomé lattice model in the imaginary field. It is interesting to compare the result Eq. (65) with Eq. (1) of Ref. Wang and Wu (2007), which is the solution of the dimer model on the Kagomé lattice and can also be derived by mapping to the odd free-fermion model. Both formulas take a very simple logarithmic form. Ref. Wang and Wu (2007) stated that the novel and unique expression “points to the special role played by the Kagomé lattice”. Here we provide new evidence for this statement.
When $J_{1}=J_{2}=J_{3}>0$, the Kagomé lattice Ising model is geometrically frustrated. The residual entropy in the imaginary field is easily calculated from Eq. (65) by taking the zero temperature limit
@@ IV Summary
All solutions in this paper are achieved by mapping the Ising spin system into a free-fermion model. The mapping procedure is straightforward and should be easy to extend to other lattice systems. Applying the mapping method to the Ising model in a physical (real) and non-zero magnetic field would be interesting. Furthermore, the free-fermion formulation may be suitable for other statistical lattice problems, e.g., the dimer covering problem. Our work has provided new insights into the mathematically exact results of spin systems, showcasing the elegance and power of the free-fermion formulation in condensed matter and statistical physics.
@@ Acknowledgements.
This work was supported by Guangdong Provincial Quantum Science Strategic Initiative (Grants No. GDZX2203001 and No. GDZX2403001), National Natural Science Foundation of China (Grants No. 12474489 and No. 12474228), Research Funding for Outbound Postdoctoral Fellows in Shenzhen (Grant No. SZRCXM2401006), Shenzhen Fundamental Research Program (Grant No. JCYJ20240813153139050), and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302300).